OpenStudy (anonymous):

Peanuts worth $4/kg were mixed with almonds worth $6/kg to produce 12 kg of nuts worth $62. A. Represent the unknowns above using the variables x and y. Show all your work. B. Write an equation that models the relationships between te weights of the nuts in the mixture. Show all your work. C. Write an equation that models the values of the nuts in the mixture. Show all your work. D. Using the two equations above, write and solve the systems to find x and y. Show all your work.

OpenStudy (anonymous):

a=almonds p=peanuts 4p+6a=(62) p+a=12 p=12-a 4(12-a) + 6a =62 48-4a+6a=62 2a=14 a=7kg p=5kg

OpenStudy (anonymous):

Thanks!!